• An isolated macroscopic system at temperature 300K absorbs a photon from the visible part of spectrum (say Wavelength=500nm) obtain a value for the relative increase ∆w/w in The number of accessible states of the system ?
1
Expert's answer
2020-11-30T14:59:53-0500

As per the given question,

Temperature (T)=300K

Wavelength "(\lambda)=500nm"

We know that "S=K_b\ln W"

Now, taking the differentiation,

"ds=K_b\frac{1}{W}dW"


"\frac{\Delta W}{W}=\frac{ \Delta S}{K_b}"

But we know that,

"\Delta S = \frac{Q}{T}=\frac{hc}{\lambda T}"

"={\frac{6.6\times 10^{-34}\times 3\times 10^8}{ 300K\times 500\times 10^{-9}}}"


"=1.326\times 10^{-21} J/K"

Hence,

"\frac{\Delta W}{W}=\frac{ \Delta S}{K_b}"


"=\frac{ 1.325\times 10^{-21}J/K}{1.38\times 10^{-23} J / K}"


"=96"